OpenStudy (anonymous):

need verification: Find a unit vector with the same direction as <4, -4, 4>

OpenStudy (anonymous):

i get \[4/\sqrt(48)i - 4/\sqrt(48)j + 4/\sqrt(48)k\]

OpenStudy (ash2326):

we have vector a=4i-4j+4k it's absolute value is squareroot(4^2+(-4)^2+4^2)=root(48)=root(16*3)=4 root 3

OpenStudy (ash2326):

yeah you're right

OpenStudy (jamesj):

So, yes, the expression simplifies to \[ \frac{1}{\sqrt{3}} ( \hat{i} - \hat{j} + \hat{k}) \]

OpenStudy (amistre64):

v/|v| is a unit if i recall it right

OpenStudy (anonymous):

i'm just reviewing for my final on thursday and getting some practice in

OpenStudy (anonymous):

making sure i understand different things